TMS Meeting Minutes, April 29, 2021

Attendees: Andrew, Breann, Heejoo, John, Matthieu, Olga, Trenton, Yang (Zoom Meeting)

Fibers Unplugged
  • Fibers were removed for eTalon to test the system today (by the German team)
  • Etalon should be able to finish the test within a few days (shouldn't take more than a week)
  • After that, the all-fiber connections will be rollback. They are well labeled.
Thorlabs Collimator Testing
  • Heejoo planning to test this once he goes up on the mountain
  • Doesn't have the hardware at Optical Sciences to test at 1550 nm
  • Planning to use a Shack Hartmann to test the system at 633 nm (as it is a broadband design)
  • eTalon does test their collimators before install with a simple spot on a wall (we can do better than this, obviously)
Binocular TMS Code
  • Yang has been working on this
  • Plans to have a draft version finished by tomorrow
  • Proposes scheduling a test next week or so during the daytime
  • Matthieu has some question as to whether it is necessary to get TMS running on both sides
  • Yang concurs with Andrew that this is no major activity, which will not cause any major overhead
  • Nice thing about this code is that it runs unobtrusively in the background
  • Yang's next steps are to put this TMS system into a database and make data accessible through DMS
Daytime Testing Plans for Next Week
  • Andrew wants to see basic sanity checks for driving M1 through standard motions to make sure we are getting the right measurements
  • Andrew requests sending a test plan, perhaps a plan produced by Yang (basically, does the software operate on both sides simultaneously)
  • Andrew to start an email chain to discuss what needs to be accomplished during this test (sending out later today)
Solid Retro
  • Heejoo sent the solid retro to Pasadena today (should take 7-10 days)
  • Tracking Number: 9505 5109 6021 1119 6485 06
Summer Break TMS Upgrade Plans
  • Potential to install the MPO fiber switch to increase the number of channels (including secondary mirror or LBTI)
  • Andrew suggests that we need to demonstrate that the fiber splitter box is a viable techonology for our system (almost none of the channels had any useful energy in them when put in use at LBT)
  • Andrew taking action item to develop the needed functionality for Gregorian mode and the ability to switch between Gregorian and prime
  • Andrew would also like to look into troubleshooting the one or two dead channels
Commissioning Plan
  • Andrew will commence the TMS commissioning plan via email
  • Heejoo wonders if we need to do this commissioning process every summer or no
  • What we are offering now with TMS is quite new and we want to offer help and not harm to the science operations at LBT
  • When things are reconfigured, or mirrors are off and on, etc., there will be some need for a soft commissioning plan (easier to test, setup)
  • Perhpas a series of system requirements that quantify proper functionality of TMS
  • However, for this particular launch, we want to have a thorough plan, kind of like a CDR, which can provide justification for the observatory to provide this system for science operations

Andrew will also send out the Karmann paper to the TMS group

Andrew is still planning to check Heejoo's SA work with calculating.

-- TrentonBrendel - 29 Apr 2021
Topic revision: r2 - 30 Apr 2021, HeejoChoi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback